For her daughter's university education, Carla Hackl has invested an inheritance in a fund paying 9.2% compounded quarterly. If ordinary annuity payments of $5750.00 per month are to be made out of the fund for 6 years and the annuity begins 7.75 years from now, how much was the inheritance?
Answer #1

C D EAR 9.52% Monthly rate Period Periodic Payment PV at t - 7.75 0.76% 72 5750 317849.88 5 9 2.30% 10 11 Quarterly rate Peri

EAR =(1+9.2%/4)^4-1 Inmt Monthly rate Period Periodic Payment PV at t - 7.75 =(1+C2)^(1/12)-1 =6*12 5750 =PV(C4,C5,-C6) 7 10